I need your help with this digit problem: The sum of the digits of a two-digit number is 12. The value of the number is two more than 11 times tens digit. Find the number.
Answer by Theo(13342)

I need your help with this digit problem: The sum of the digits of a two-digit number is 12. The value of the number is two more than 11 times tens digit. Find the number.
-----
x = first digit
y = second digit
-----
The sum of the digits of a two-digit number is 12.
-----
this means that x + y = 12
x is the tens digit.
y is the units digit.
-----
The value of the number is two more than 11 times tens digit.
-----
10 * x + y is the value of the number.
11 * x + 2 is two more than the value of the tens digit * 11.
Your equation here is:
10*x + y = 11*x + 2
This says that the value of your number (10*x + y) is 2 more than 11 times the tens digit (11*x + 2)
-----
You have 2 equations that need to be solved simultaneously.
They are:
-----
x + y = 12
and
10*x + y = 11*x + 2
-----
You can solve for y in your second equation of 10*x + y = 11*x + 2 to get:
y = x + 2
-----
You can substitute for y in your first equation to get:
x + (x + 2) = 12
-----
Remove parentheses to get:
2x + 2 = 12
Solve for x to get:
x = 5
-----
If x = 5, then x + y = 12 becomes:
5 + y = 12
Solve for y to get:
y = 7
-----
You have:
x = 5
y = 7
-----
Sum of the digits is 12 which is good.
-----
The number is xy = 57
-----
The value of the number is 2 more than 11 times the tens digit.
11 * the tens digit is 11 * 5 = 55
55 + 2 = 57
This is true confirming the values for x and y are good.
-----
Your answer is:
Number is 57
-----